# Nonstrange and strange quark matter at finite temperature within modified NJL model and protoquark stars

### v1

@@ Abstract
We extend the modified Nambu–Jona-Lasinio (NJL) model—incorporating exchange interactions via the Fierz transformation—to finite temperatures in both two- and three-flavor scenarios, and investigate the properties of protoquark stars in $\beta$-equilibrium. Our results show that increasing the strength of exchange interactions, characterized by the parameter $\alpha$, changes the chiral phase transition from first-order to crossover. We examine the effects of finite temperature, lepton fraction, and exchange interactions on the equation of state (EOS), finding that the EOS in the crossover regime is significantly stiffer than in the first-order case. In certain regions of parameter space within the crossover scenario, strange quark matter can support the existence of twin star configurations.
@@ I Introduction
Neutron stars are the remnants of supernova explosions. They typically have masses in the range of $1-2\,M_{\odot}$, radii on the order of $\sim 10$ km, and are born with temperatures as high as $10^{11}$ K. Within a few days, they cool to approximately $10^{10}$ K by emitting neutrinos. In stellar modeling, the internal structure of a neutron star is determined by the assumed equation of state (EOS), which remains uncertain due to the incomplete understanding of the strong interaction under extreme conditions. At present, extensive studies have been conducted on proto-neutron stars with or without quarks at finite temperature Prakash et al. (1997); Reddy et al. (1998); Steiner et al. (2000); Pons et al. (2001); Steiner et al. (2001); Nicotra et al. (2006); Burgio & Plumari (2008); Yasutake & Kashiwa (2009); Shao (2011); Chen et al. (2012); Menezes et al. (2015); Raithel et al. (2019).
@@
According to the Bodmer-Witten hypothesis Bodmer (1971); Witten (1984), compact stars can be composed entirely of deconfined quarks, forming self-bound quark stars. The composition of these self-bound stars hinges on a fundamental question—the true ground state of baryonic matter at zero pressure and temperature. After decades of theoretical speculation, both strange quark stars, composed of strange quark matter Chakrabarty et al. (1989); Dey et al. (1998); Chakrabarty (1993); Buballa & Oertel (1999); Peng et al. (1999); Wen et al. (2005); Li et al. (2010); Zhou et al. (2018); Li et al. (2018); Xia et al. (2019); Peng et al. (2000); Bai & Liu (2021); Xia et al. (2021); Yuan et al. (2022); Zhou et al. (2024); Zhang et al. (2024), and up-down quark stars, containing only up and down quark matter Holdom et al. (2018); Zhao et al. (2019); Ren & Zhang (2020); Zhang (2020); Zhang & Mann (2021); Restrepo et al. (2023); Su et al. (2024); Yang et al. (2025), have emerged as viable alternative models to neutron stars. Potential candidates for quark stars include compact objects with gravitational masses $\lesssim 1\,M_{\odot}$, such as PSR J1645–0317 (PSR B1642–03) Taylor et al. (1993), PSR J1830–1059 (PSR B1828–11) Taylor et al. (1993), 4U 1728–34 Li et al. (1999), and the X-ray bursters GRO J1744–28 Cheng et al. (1998) and SAX J1808.4–3658 Li et al. (1999). A recent analysis of the supernova remnant HESS J1731-347 suggests that its central compact object has a remarkably low mass of $M=0.77^{+0.20}_{-0.17}\,M_{\odot}$ and a radius of $R=10.4^{+0.86}_{-0.78}$  km Doroshenko et al. (2022), reigniting interest in the possible existence of self-bound quark stars. While observations of isolated compact stars provide constraints on the EOS of cold dense matter, many dynamical phenomena–such as core-collapse supernovae, the formation and cooling of proto-compact stars, and compact star mergers–depend on the EOS at finite temperatures. Thus, further investigation into the thermal effects on quark matter and the evolution of proto-quark stars is both crucial and intellectually stimulating Menezes et al. (2006); Dexheimer et al. (2013, 2014); Chu et al. (2017, 2019).
@@
Due to the nonperturbative challenges at finite baryon densities, effective models that incorporate key features of QCD offer a practical alternative. For example, Ref. Lopes et al. (2021) employed the modified MIT bag model to study the quark matter at finite temperature and hot stars. To effectively capture the dynamics of spontaneous chiral symmetry breaking in QCD, the Nambu–Jona-Lasinio (NJL) model 1992RvMP…64..649K ; Hatsuda & Kunihiro (1994); 2005PhR…407..205B serves as a particularly suitable choice, which successfully reproduces the spectrum of low-lying mesons 2005PhR…407..205B ; 1992RvMP…64..649K ; Hatsuda & Kunihiro (1994), and has been widely applied and extended to describe quark matter in compact star physics (see, e.g., Refs. Huang et al. (2003); Blaschke et al. (2005, 2005); Wang et al. (2019); Minamikawa et al. (2021); Xia (2024); Gholami et al. (2025, 2024) as incomplete lists). Within this framework, Ref. Chu et al. (2017) investigates the role of isovector interactions and the behavior of the symmetry free energy of quark matter at finite temperature. In Ref. Menezes et al. (2006), the authors compare the predictions of the NJL model and the MIT bag model for quark matter at finite temperature and the properties of quark stars.
@@
To improve the description of strongly interacting matter at high baryon chemical potentials, a modified version of the NJL model has been proposed in Ref. Wang et al. (2019). This extended NJL model incorporates both the original NJL Lagrangian and its Fierz-transformed counterpart, with their respective contributions weighted by the $1-\alpha$ and $\alpha$ parameters. This modified NJL model has been successfully applied in various studies, including the study of the critical endpoint in the QCD phase diagram Yang et al. (2019); Yu et al. (2020), color superconductivity Su et al. (2020); Yuan et al. (2023); Yuan & Li (2024); Yang et al. (2025), as well as the properties of cold quark matter and quark stars Wang et al. (2020); Yuan et al. (2022); Su et al. (2024). In this work, as an extension to previous studies, we first employ this modified NJL-type model to describe two-flavor and three-flavor quark matter at finite temperature, both with and without trapped neutrinos. We then systematically investigate the impact of exchange interactions on the properties of chiral phase transition and explore their implications for both strange and nonstrange protoquark stars.
@@
This paper is organized as follows. In Sec. II, we will present the modified NJL model employed in this work to describe the nonstrange and strange quark matter at finite temperature, including the Fierz transformed interactions. Section III and Sec. V discuss the results on quark matter EOS at finite temperature and proto-quark stars. Our results are summarized in Sec. VI.
@@ II The modified NJL models
Here, $\mathcal{L}_{0}$ denotes the relativistic free field, $\mathcal{L}_{0}=\bar{\psi}\left(i\gamma^{\mu}\partial_{\mu}-m+\mu\gamma^{0}\right)\psi$, which describes the propagation of non-interacting fermions. $G$ is the four-fermion interaction coupling constant. $\psi$ is the quark field operator with color, flavor, and Dirac indices. $\mu$ is the flavor-dependent quark chemical potential. The quark mass matrix $m=\mathrm{diag}(m_{u},m_{d})$ is diagonal in flavor space and includes the small current quark masses, thereby introducing a slight explicit breaking of chiral symmetry.
For further considering the effect of a rearrangement of fermion field operators, we apply the Fierz transformation to the interaction terms in the NJL models, as discussed in Refs. Wang et al. (2019); Yang et al. (2019); Yu et al. (2020); Su et al. (2020); Yuan et al. (2023); Yuan & Li (2024); Wang et al. (2020); Yuan et al. (2022); Su et al. (2024). As a purely technical device to examine the exchange channels influence that occur in quartic products at the same space-time point 1992RvMP…64..649K ; 2005PhR…407..205B , the Fierz identity of the four-fermion interactions in the two-flavor NJL model is
@@
$\left.+\left(\bar{\psi}\sigma^{\mu\nu}\psi\right)^{2}-\left(\bar{\psi}\sigma^{\mu\nu}\tau\psi\right)^{2}\right]\ .$
We note that we only consider the contribution of color singlet terms for simplicity. One can see that, through the Fierz transformation, all exchange interaction channels of the original Lagrangian are released. In Eq. (2), the Fierz transformed Lagrangian contains not only the scalar and pseudoscalar interactions, but also vector and axialvector interaction channels.
@@
When $\alpha=0$, the effective Lagrangian reduces to the standard NJL model. Unlike the conventional approach of artificially adding the vector interaction to the standard NJL Lagrangian when studying systems at finite chemical potential, the modified Lagrangian in Eq. (3) introduces vector interactions self-consistently Wang et al. (2019).
Under the mean-field approximation, the scalar and vector interactions in the Fierz-transformed channels in Eq. (2) also contribute at finite densities, thereby modifying the effective quark mass $M$ and the effective quark chemical $\tilde{\mu}$. The corresponding mass gap equation and the effective chemical potential for each quark are given as follows:
@@
In finite-temperature field theory Kapusta and Gale (2011), the linearization of $\mathcal{L}_{\rm eff}$ in the vicinity of the expectation values and the application of Matsubara formalism yield the thermodynamical potential per volume of quark matter as follows:
@@
with
@@
which is the Fermi-Dirac distribution with $E_{f}=\sqrt{p^{2}+M_{f}^{2}}$.
II.0.2 (2+1)-flavor modified NJL model
The Lagrangian density of the (2+1)-flavor NJL model is given by
@@
where $\mathcal{L}_{\sigma}^{4}$ and $\mathcal{L}_{\sigma}^{6}$ represent the four-fermion and six-fermion interaction terms, respectively. These interaction terms are expressed as
$\mathcal{L}_{\sigma}^{4}$ $=\sum_{i=0}^{8}G\left[\left(\bar{\psi}\lambda_{i}\psi\right)^{2}+\left(\bar{\psi}i\gamma^{5}\lambda_{i}\psi\right)^{2}\right]\ ,$ (10)
$\mathcal{L}_{\sigma}^{6}$ $=-K\left(\operatorname{det}\left[\bar{\psi}\left(1+\gamma^{5}\right)\psi\right]+\operatorname{det}\left[\bar{\psi}\left(1-\gamma^{5}\right)\psi\right]\right)\ .$
Here, $G$ and $K$ denote the coupling constants for the four-fermion and six-fermion interactions, respectively. The matrices $\lambda_{i}\;(i=1\rightarrow 8)$ correspond to the Gell-Mann matrices in flavor space, while $\lambda_{0}=\sqrt{2/3}\;I_{0}$, with $I_{0}$ being the identity matrix. Applying the Fierz transformation to the four-fermion scalar and pseudoscalar interaction terms, considering only the contributions from color-singlet channels, yields
@@
In the mean-field approximation, the mass gap equations and the effective chemical potential $\mu_{f}^{*}$ for a given flavor $f=i,j,k$ are given by
@@
The grand canonical potential density of the interacting strange quark matter system is
@@
$-\frac{\alpha G}{3}\left(\rho_{u}+\rho_{d}+\rho_{s}\right)^{2}-4K\sigma_{u}\sigma_{d}\sigma_{s}+\Omega_{0}\ .$
Then all thermodynamic quantities of interest can be calculated. The pressure is
@@
Here, we choose the constant in Eq. (15) such that $P$ and $\epsilon$ vanish in vacuum, i.e., we choose $\Omega_{M}(0,0)=0$.
@@ III Quark matter in $\beta$-equilibrium
A newborn neutron star forms in the aftermath of a successful supernova explosion, when the stellar remnant becomes gravitationally decoupled from the expanding ejecta and reaches core temperatures as high as $\sim 50\;\text{MeV}$ Lattimer & Prakash (2004); Weber (2005). In this work, we consider two distinct scenarios: an early stage in which neutrinos are trapped within the stellar interior, and a later stage after deleptonization, where neutrinos have escaped and their chemical potential vanishes. Although an isentropic description is generally considered more realistic for modeling proto-quark stars, we adopt an isothermal approach, as made in previous work Santos & Menezes (2004); Yasutake & Kashiwa (2009); Lu et al. (2019); Lopes et al. (2021); Pal & Chaudhuri (2024), to enable direct comparisons across different temperature cases.
@@
One has to require also the charge-neutrality of the quark matter,
@@
The lepton has no interactions, so the grand canonical potential density is the free grand canonical potential density. The lepton densities are given by
@@
The EOS of dense warm matter is obtained by including both the contribution of quarks and leptons, namely the total energy density, pressure and entropy density are given by
@@
Figure 1: Upper panel: The effective quark masses of u,d,su,\ d,\ s quarks versus quark chemical potential $\mu$ both for strange quark matter and nonstrange quark matter in the absence of neutrinos ($Y_{l}=0.0$), under various values of $\alpha$ and temperature $T$. The results for several representative parameter sets are shown: ($\alpha=0.0,\ T=0.04$ GeV), ($\alpha=0.0,\ T=0.00$ GeV), ($\alpha=0.5,\ T=0.00$ GeV), ($\alpha=0.8,\ T=0.00$ GeV), ($\alpha=0.9,\ T=0.00$ GeV). Middle panel: The corresponding results for quark number densities of u,d,su,\ d,\ s quarks as functions of $\mu$. Lower panel: The electron chemical potential $\mu_{e}$ as a function of $\mu$.
Figure 2: The effective quark masses $M_{u}$, $M_{d}$, and $M_{s}$, the electron chemical potential $\mu_{e}$, and the $d$ quark chemical potential $\mu_{d}$ are shown as functions of the $u$ quark chemical potential in the neutrino-trapped scenario, for a representative case with $\alpha=0.5$, $Y_{l}=0.1$, and $T=0.04$ GeV. Results are presented for both strange and nonstrange quark matter.
@@
Before doing the calculation, we should fit the NJL model parameters. Since the NJL model is not renormalizable, a momentum cutoff $\Lambda$ should be introduced to regulate the ultraviolet divergences. Here, we use the three momentum cutoff. After defining the new coupling constant $G^{\prime}$, and keeping the expression of gap equation the same as the widely used one in Eqs. (4) and (14), at zero temperature and chemical potential, apart from $\alpha$, the fixing of the model parameters is the same with the original version of the NJL model Hatsuda & Kunihiro (1994). Similar to the procedure in Ref. Hatsuda & Kunihiro (1994), after fixing the masses of the up and down quarks by equal values, the other parameters $m_{s},\Lambda_{\mathrm{UV}},G^{\prime}$, $K$ are chosen to reproduce the experimental data of the pion decay constant and pion mass. The employed model parameters are shown in Table 1.
@@
Two-flavor $5.0$ / $653.0$ $4.930\times 10^{-6}$ /
$(2+1)$-flavor $5.5$ 104.7 602.3 $5.058\times 10^{-6}$ $1.559\times 10^{-13}$
To obtain the EOS, the following set of equations must be solved self-consistently. For neutrino-free matter, the relevant system consists of the gap equations of the NJL model in Eqs. (4) and (14), along with the conditions for $\beta$-equilibrium in Eq. (21) and charge neutrality in Eq. (22). These constraints reduce the number of independent chemical potentials to one. We choose the up quark chemical potential $\mu_{u}$ as the input quantity, with the remaining chemical potentials, $\mu_{d}$, $\mu_{s}$ and $\mu_{e}$ determined as functions of $\mu_{u}$. In the case of neutrino-trapped matter, the system includes the same NJL gap equations, together with the modified $\beta$ equilibrium condition in Eq. (24), charge neutrality in Eq. (22), and a fixed lepton fraction in Eq. (25). Again, only one independent chemical potential remains.

### v2

@@ Abstract
We extend the modified Nambu–Jona-Lasinio (NJL) model–incorporating exchange interactions via the Fierz transformation–to finite temperatures in both two- and three-flavor scenarios, and investigate the properties of protoquark stars in $\beta$-equilibrium. Our results show that increasing the strength of exchange interactions, characterized by the parameter $\alpha$, changes the chiral phase transition from first-order to crossover. We examine the effects of finite temperature, lepton fraction, and exchange interactions on the equation of state (EOS). We find that, in the crossover regime, the EOS is significantly stiffer than in the first-order case due to the substantial contribution of repulsive interactions in the exchange channels, while it remains relatively insensitive to variations in temperature and lepton fraction. Imposing the thermodynamic consistency, which requires the minimum of $f/\rho_{B}$ occurs at zero pressure, further constrains the minimum value of vacuum pressure.
@@ I Introduction
Neutron stars are the remnants of supernova explosions. They typically have masses in the range of $1-2\,M_{\odot}$, radii on the order of $\sim 10$ km, and are born with temperatures as high as $10^{11}$ K. Within a few days, they cool to approximately $10^{10}$ K by emitting neutrinos. In stellar modeling, the internal structure of a neutron star is determined by the assumed equation of state (EOS), which remains uncertain due to the incomplete understanding of the strong interaction under extreme conditions. At present, extensive studies have been conducted on proto-neutron stars with or without quark cores at finite temperature Prakash et al. (1997); Reddy et al. (1998); Steiner et al. (2000); Pons et al. (2001); Steiner et al. (2001); Nicotra et al. (2006); Burgio & Plumari (2008); Yasutake & Kashiwa (2009); Shao (2011); Chen et al. (2012); Menezes et al. (2015); Raithel et al. (2019).
@@
According to the Bodmer-Witten hypothesis Bodmer (1971); Witten (1984), compact stars can be composed entirely of deconfined quarks, forming self-bound quark stars. The composition of these self-bound stars hinges on a fundamental question: what is the true ground state of baryonic matter at zero pressure and temperature? After decades of theoretical speculation, both strange quark stars, composed of strange quark matter Chakrabarty et al. (1989); Dey et al. (1998); Chakrabarty (1993); Buballa & Oertel (1999); Peng et al. (1999); Wen et al. (2005); Li et al. (2010); Zhou et al. (2018); Li et al. (2018); Xia et al. (2019); Peng et al. (2000); Bai & Liu (2021); Xia et al. (2021); Yuan et al. (2022); Zhou et al. (2024); Zhang et al. (2024), and up-down quark stars, containing only up and down quark matter Holdom et al. (2018); Zhao et al. (2019); Zhang (2020); Zhang & Mann (2021); Restrepo et al. (2023); Su et al. (2024); Yang et al. (2025), have emerged as viable alternative models to neutron stars. Besides, self-bound stars composed of strangeon matter have also been proposed as candidates Xu (2003); Lai & Xu (2009); Zhang et al. (2023); Yuan et al. (2025). Potential candidates for quark stars include compact objects with gravitational masses $\lesssim 1\,M_{\odot}$, such as PSR J1645-0317 (PSR B1642-03) Taylor et al. (1993), PSR J1830-1059 (PSR B1828-11) Taylor et al. (1993), 4U 1728-34 Li et al. (1999), and the X-ray bursters GRO J1744-28 Cheng et al. (1998) and SAX J1808.4-3658 Li et al. (1999). A recent analysis of the supernova remnant HESS J1731-347 suggests that its central compact object has a remarkably low mass of $M=0.77^{+0.20}_{-0.17}\,M_{\odot}$ and a radius of $R=10.4^{+0.86}_{-0.78}$  km Doroshenko et al. (2022), reigniting interest in the possible existence of quark stars. While observations of isolated compact stars provide valuable constraints on the EOS of cold dense matter, many dynamical astrophysical phenomena–such as core-collapse supernovae, the formation and cooling of proto-compact stars, and compact star mergers–depend on the EOS at finite temperatures. To account for thermal effects, an ideal-fluid thermal component is commonly introduced into an arbitrary cold EOS Janka et al. (1993). In this prescription, the thermal contribution is parameterized by a simple adiabatic index as $P_{th}=\epsilon_{th}(\Gamma_{th}-1)$, where $P_{th}$ and $\epsilon_{th}$ are the thermal pressure and energy density, and $\Gamma_{th}$ is the adiabatic index, which is assumed to be constant. Although this treatment offers computational simplicity, it neglects the effects of degeneracy on the thermal pressure and lacks explicit dependence on temperature and entropy. While several nuclear-matter EOSs have been developed to incorporate finite-temperature effects self-consistently Lattimer & Swesty (1991); Shen et al. (1998); Hempel & Schaffner-Bielich (2010); Zhu et al. (2025), only a limited number of them exist, and even fewer account for phase transitions or exotic degrees of freedom. Therefore, further investigation of thermal effects in quark matter and the evolution of protoquark stars within effective field theory frameworks remains both essential and intellectually compelling Menezes et al. (2006); Dexheimer et al. (2013, 2014); Chu et al. (2017, 2019).
@@
Due to the nonperturbative challenges at finite baryon densities, effective models that incorporate key features of QCD offer a practical alternative. For example, Ref. Lopes et al. (2021) employed the modified MIT bag model to study the quark matter at finite temperature and hot stars. To effectively capture the dynamics of spontaneous chiral symmetry breaking in QCD, the Nambu-Jona-Lasinio (NJL) model 1992RvMP…64..649K ; 2005PhR…407..205B serves as a particularly suitable choice, which successfully reproduces the spectrum of low-lying mesons 2005PhR…407..205B ; 1992RvMP…64..649K , and has been widely applied and extended to describe quark matter in compact star physics (see, e.g., Refs. Huang et al. (2003); Blaschke et al. (2005, 2005); Wang et al. (2019); Minamikawa et al. (2021); Xia (2024); Gholami et al. (2025, 2024) as incomplete lists). Within this framework, Ref. Chu et al. (2017) investigates the role of isovector interactions and the behavior of the symmetry free energy of strange quark matter at finite temperature. In Ref. Menezes et al. (2006), the authors compare the predictions derived from the standard NJL model containing scalar interactions with those of the MIT bag model for strange quark matter EOS at finite temperature, as well as the properties of strange quark stars.
@@
To improve the description of strongly interacting matter at high baryon chemical potentials, a modified version of the NJL model has been proposed in Ref. Wang et al. (2019). This extended NJL model incorporates both the original NJL Lagrangian and its Fierz-transformed counterpart, with their respective contributions weighted by the parameters $1-\alpha$ and $\alpha$. The parameter $\alpha$ is treated as a free parameter, allowed to vary between 0 to $1$, enabling the model to accommodate constraints at finite baryon density. In this sense, the present formulation generalizes the NJL-type Lagrangian, of which the specific choice $\alpha=0.5$ introduced in Ref. 1992RvMP…64..649K represents a special case. This modified NJL model has been successfully applied in various studies, including the study of the chiral chemical potential $\mu_{5}$ and its physical effects in two-flavor quark matter system Yang et al. (2019); Wang et al. (2021), investigations of the critical endpoint in the QCD phase diagram within the three-flavor version of the model and its comparison with the two-flavor case Yu et al. (2020), as well as the the absolute stability and phase structure of two-flavor color superconductivity Su et al. (2020); Yuan et al. (2023); Yuan & Li (2024); Yang et al. (2025). Compact stars, serving as natural laboratories to explore the nature of dense matter, have strongly motivated studies of nonstrange and strange quark stars Wang et al. (2020); Yuan et al. (2022); Su et al. (2024), as well as neutron stars with quark-matter cores Yuan et al. (2024, 2025), within the framework of the modified NJL model with exchange interactions parameterized by $\alpha$ to describe the cold quark matter phase. To our knowledge, existing studies within the NJL-like model have primarily focused on cold quark matter Wang et al. (2020); Yuan et al. (2022); Su et al. (2024) or thermal three-flavor quark matter systems Menezes et al. (2006); Chu et al. (2017). However, the case of nonstrange quark matter and its systematic comparison with strange quark matter at finite temperature–particularly regarding the properties of nonstrange and strange protoquark stars–has not yet been explored in detail. Extending these earlier efforts, in this work, we employ this modified NJL-type model to describe two- and three-flavor quark matter with and without trapped neutrinos at finite temperature. We then systematically investigate the impact of exchange interactions on the characteristics of chiral phase transition and explore their implications for both strange and nonstrange protoquark stars.
@@
This paper is organized as follows. In Sec. II, we will present the modified NJL model employed in this work to describe the nonstrange and strange quark matter at finite temperature, including the Fierz transformed interactions. Sections IIIIV and Sec. V discuss the results on quark matter EOS with and without trapped neutrinos at finite temperature and protoquark stars. Our results are summarized in Sec. VI.
@@ II The modified NJL models
Here, $\mathcal{L}_{0}$ denotes the relativistic free field, which describes the propagation of non-interacting fermions. $G$ is the four-fermion interaction coupling constant. $\psi$ is the quark field operator with color, flavor, and Dirac indices. $\mu$ is the flavor-dependent quark chemical potential. The quark mass matrix $m=\mathrm{diag}(m_{u},m_{d})$ is diagonal in flavor space and includes the small current quark masses, thereby introducing a slight explicit breaking of chiral symmetry.
For further considering the effect of a rearrangement of fermion field operators, we apply the Fierz transformation to the interaction terms in the NJL models Wang et al. (2019). As a purely technical device to examine the exchange channels influence that occur in quartic products at the same space-time point 1992RvMP…64..649K ; 2005PhR…407..205B , the Fierz identity of the four-fermion interactions in the two-flavor NJL model is
@@
$\left.+\left(\bar{\psi}\sigma^{\mu\nu}\psi\right)^{2}-\left(\bar{\psi}\sigma^{\mu\nu}\tau\psi\right)^{2}\right]\ ,$
with only considering the contribution of color singlet terms for simplicity. One can see that, through the Fierz transformation, all exchange interaction channels of the original Lagrangian are released. In Eq. (2), the Fierz transformed Lagrangian contains not only the scalar and pseudoscalar interactions, but also vector and axial-vector interaction channels.
@@
When $\alpha=0$, the effective Lagrangian reduces to the standard NJL model. Unlike the conventional approach of artificially adding the vector interaction to the standard NJL Lagrangian for systems at finite chemical potential, the modified Lagrangian in Eq. (3) introduces vector interactions self-consistently Wang et al. (2019).
Under the mean-field approximation, the scalar and vector interactions in the Fierz-transformed channels in Eq. (2) also contribute at finite densities, thereby modifying the constituent quark mass $M$ and the effective quark chemical $\tilde{\mu}$. The corresponding mass gap equation and the effective chemical potential for each quark are given as follows:
@@
In finite-temperature field theory Kapusta and Gale (2011), the linearization of $\mathcal{L}_{\rm eff}$ in the vicinity of the expectation values and the application of Matsubara formalism yield the thermodynamical potential per volume of quark matter expressed as:
@@
where $n_{f}$ and $\bar{n}_{f}$ are Fermi occupation numbers of quarks and antiquarks, respectively,
@@
with $E_{f}=\sqrt{p^{2}+M_{f}^{2}}$.
II.0.2 three-flavor modified NJL model
The Lagrangian density of the three-flavor NJL model is given by
@@
$\mathcal{L}_{\sigma}^{4}=\sum_{i=0}^{8}G\left[\left(\bar{\psi}\lambda_{i}\psi\right)^{2}+\left(\bar{\psi}i\gamma^{5}\lambda_{i}\psi\right)^{2}\right]\ ,$
$\mathcal{L}_{\sigma}^{6}=-K\left(\operatorname{det}\left[\bar{\psi}\left(1+\gamma^{5}\right)\psi\right]+\operatorname{det}\left[\bar{\psi}\left(1-\gamma^{5}\right)\psi\right]\right)\ ,$
where $\mathcal{L}_{\sigma}^{4}$ and $\mathcal{L}_{\sigma}^{6}$ represent the four-fermion and six-fermion interaction terms, respectively. $G$ and $K$ denote the coupling constants for the four-fermion and six-fermion interactions. The matrices $\lambda_{i}\;(i=1\rightarrow 8)$ correspond to the Gell-Mann matrices in flavor space. $\lambda_{0}=\sqrt{2/3}\;I_{0}$, with $I_{0}$ being the identity matrix. Applying the Fierz transformation to the four-fermion scalar and pseudoscalar interaction terms, considering only the contributions from color-singlet channels, yields
@@
In the mean-field approximation, the mass gap equations and the effective chemical potential $\mu_{f}^{*}$ for a given flavor $(f=i,j,k)$ are expressed as
@@
From the grand canonical potential density of the interacting strange quark matter system
@@
$-\frac{\alpha G}{3}\left(\rho_{u}+\rho_{d}+\rho_{s}\right)^{2}-4K\sigma_{u}\sigma_{d}\sigma_{s}+\Omega_{0}\ ,$
the pressure can be calculated as follows:
@@
Here, we choose the constant $\Omega_{0}$ in Eqs. (5) and (14) such that $P$ and $\epsilon$ vanish in the vacuum, i.e., $\Omega_{M}(0,0)=0$.
@@ III Quark matter in $\beta$-equilibrium
A newborn neutron star forms in the aftermath of a successful supernova explosion, when the stellar remnant becomes gravitationally decoupled from the expanding ejecta and core temperatures reaches as high as $\sim 50\;\text{MeV}$ Lattimer & Prakash (2004); Weber (2005). In this work, we consider two distinct scenarios: an early stage in which neutrinos are trapped within the stellar interior, and a later stage after deleptonization, where neutrinos have escaped and their chemical potential vanishes. Although an isentropic description is generally considered more realistic for modeling protoquark stars, we adopt an isothermal approach to enable direct comparisons across different temperature cases.
@@
To describe stellar matter, the charge-neutrality of the quark matter should also be required,
@@
The lepton has no interactions, so the grand canonical potential density is the free grand canonical potential density. The lepton number densities are given by
@@
The EOS of dense warm matter is obtained by including both the contribution of quarks and leptons, namely the total energy density, pressure and entropy density are
@@
The free energy density of the system is
$f=\epsilon-TS~.$ (27)
In the case of isothermal scenario considered in the present work, where the temperature remains constant, the pressure can be expressed as
$P=\rho_{B}^{2}\frac{\partial}{\partial\rho_{B}}\left(\frac{f}{\rho_{B}}\right)_{T}~,$ (28)
which implies that the minimum of $f/\rho_{B}$ should occur at zero pressure Fisher (1967); Dexheimer et al. (2013). This condition, stemming from thermodynamic self-consistency, will be used as a criterion to further assess the EOS and to test its thermodynamic consistency in Sec. V.
@@
Before doing the calculation, we should fit the NJL model parameters. Since the NJL model is not renormalizable, here, we use the three-momentum cutoff to regulate the ultraviolet divergences. After defining the new coupling constant $G^{\prime}$, and keeping the expression of gap equation the same as the widely used one in Eqs. (4) and (13), at zero temperature and chemical potential, apart from $\alpha$, the fixing of the model parameters is the same with the original version of the NJL model Rehberg et al. (1996). For two-flavor NJL model, the $\Lambda_{\rm UV}$ and $G^{\prime}$ are determined by fitting the experimental observables of pion decay constant and pion mass Wang et al. (2019). For the three-flavor case, we adopt the parameter sets of Rehberg, Klevansky, and Hüfner (RKH) Rehberg et al. (1996). This group sets $m_{u,d}=5.5\;\text{MeV}$, and the remaining four parameters, $m_{s},\Lambda_{\mathrm{UV}},G^{\prime}$, $K$, are chosen to reproduce the experimental data of the pion decay constant $f_{\pi}=92.4\;\text{MeV}$, the pion mass $m_{\pi}=135.0\;\text{MeV}$, the kaon mass $m_{K}=497.7\;\text{MeV}$, and the $\eta^{\prime}$ meson mass $m_{\eta^{\prime}}=957.8\;\text{MeV}$.
Figure 1: Upper panel: The constituent quark masses of u,d,su,\ d,\ s quarks versus quark chemical potential $\mu$ both for strange quark matter and nonstrange quark matter in the absence of neutrinos ($Y_{l}=0.0$), under various values of $\alpha$ and temperature $T$. The results for several representative parameter sets are shown: ($\alpha=0.0$, $T=40$ MeV), ($\alpha=0.0$, $T=0$ MeV), ($\alpha=0.5$, $T=0$ MeV), ($\alpha=0.8$, $T=0$ MeV), ($\alpha=0.9$, $T=0$ MeV). Middle panel: The corresponding results for quark number densities of u,d,su,\ d,\ s quarks as functions of $\mu$. Lower panel: The electron chemical potential $\mu_{e}$ as a function of $\mu$.
Figure 2: The constituent quark masses $M_{u}$, $M_{d}$, and $M_{s}$, the electron chemical potential $\mu_{e}$, and the $d$ quark chemical potential $\mu_{d}$ are shown as functions of the $u$ quark chemical potential in the neutrino-trapped scenario, for a representative case with $\alpha=0.5$, $Y_{l}=0.1$, and $T=40$ MeV. Results are presented for both strange and nonstrange quark matter.
@@
Two-flavor Wang et al. (2019) $5.0$ / $653.0$ $4.930\times 10^{-6}$ /
Three-flavor Rehberg et al. (1996) $5.5$ 140.7 602.3 $5.058\times 10^{-6}$ $1.559\times 10^{-13}$
To obtain the EOS, the following set of equations must be solved self-consistently. For neutrino-free matter, the relevant system consists of the gap equations of the NJL model in Eq. (4) or Eq. (13), along with the conditions for $\beta$-equilibrium in Eq. (20) and charge neutrality in Eq. (21). These constraints reduce the number of independent chemical potentials to one. We choose the up quark chemical potential $\mu_{u}$ as the input quantity, with the remaining chemical potentials, $\mu_{d}$, $\mu_{s}$ and $\mu_{e}$, determined as functions of $\mu_{u}$. In the case of neutrino-trapped matter, the system includes the same NJL gap equations, together with the modified $\beta$ equilibrium condition in Eq. (23), charge neutrality in Eq. (21), and a fixed lepton fraction in Eq. (24). Again, only one independent chemical potential remains.